# One DinomalyDinomaly2 Detect Them All: A Unified Framework for Full-Spectrum Unsupervised Anomaly Detection

### v1

@@ 1 Introduction
Early image UAD approaches follow a one-class-one-model paradigm[7, 8], where separate models are trained for each object category or scene (Fig. 1(a.1)). While this strategy has proven effective in controlled scenarios, it faces significant scalability challenges in real-world applications. The storage overhead, computational complexity, and maintenance burden of managing hundreds of individual models become prohibitive in practical deployments. Recent advances have shifted toward multi-class UAD (MUAD, Fig. 1(a.2)), where a single unified model is designed to handle multiple object categories simultaneously[9, 10, 11]. Recent methods leverage a variety of techniques, including neighbor-masked attention [9], synthetic anomalies [10], vector quantization [12], diffusion model [13, 14], and state space model (Mamba) [15] to capture compact representations of normal patterns, which are critical for distinguishing anomalies. However, when intra-class normal patterns become excessively complex due to the presence of diverse categories, the underlying distribution becomes difficult to characterize, consequently degrading detection performance [9]. Yet despite this proliferation of complex designs, there is still a non-negligible performance gap between the state-of-the-art (SOTA) multi-class models and single-class models, restricting the practicability of implementing unified models, as shown in Fig. 1(d).
@@
Moreover, we present simple strategies to extend vanilla 2D Dinomaly2 to achieve SOTA results on diverse scenarios including multi-view inspection, RGB+3D point cloud, RGB+Infrared (IR), and few-shot UAD, etc (Fig. 1(b.2-4)). We further advance the field by introducing the inference-unified MUAD setting (Fig. 1(a.3)), where anomalies across mixed categories must be detected with a single threshold—a crucial step toward truly unified deployment. This setting exposes the limitations of existing methods that rely on category-specific calibration while demonstrating Dinomaly2’s robustness in this challenging setting.
@@ 2 Related Work
Unsupervised Anomaly Detection for computer vision has evolved through several paradigms, each addressing the fundamental challenge of identifying abnormal patterns without explicit anomaly supervision. Epistemic methods are based on the assumption that the networks respond differently during inference between seen input and unseen input [20]. Within this paradigm, pixel reconstruction [30, 31] methods assume that the networks trained on normal images can reconstruct anomaly-free regions well, but poorly for anomalous regions. However, such models may also incorrectly reconstruct anomalous regions if these regions resemble normal regions[8]. Therefore, feature reconstruction (distillation) [8, 32, 33, 34] is proposed to construct features of pre-trained encoders instead of raw pixels. Recent advances [14, 35] have explored diffusion models for pixel and feature reconstruction. Synthetic anomaly methods generate pre-defined defects on normal images to imitate anomalies, converting UAD into supervised classification [36] or segmentation tasks [37, 38]. For example, CutPaste [36] simulates anomalous regions by randomly pasting cropped patches of normal images, while DRAEM [37] and DeTSeg [39] construct anomalous regions using Perlin noise as the mask and another image as the additive anomaly. SimpleNet [40] introduces anomaly by injecting Gaussian noise in the pre-trained feature space. These methods heavily rely on how well the synthetic anomalies match the real anomalies, which makes it hard to generalize to different datasets. Feature statistics (or memory-based) methods [41, 7, 42] memorize all normal features (or their modeled distribution) extracted by networks pre-trained on large-scale datasets and match them with test samples during inference, e.g., PatchCore[7]. Since these methods require memorizing, processing, and matching nearly all features from training samples, they are computationally expensive in both training and inference, especially when the training set is large.
@@ 3.4 Context-Aware Recentering
A fundamental challenge in multi-class UAD arises from the context-dependent nature of anomalies: identical visual features can be normal in one context but anomalous in another. For example, a vehicle is normal on a highway but anomalous on a pedestrian walkway, whereas a pedestrian is expected on a sidewalk but becomes anomalous on a highway. This contextual ambiguity becomes particularly problematic in unified multi-class models, where the decoder may learn to reconstruct both vehicles and pedestrians in any context, undermining the anomaly detection capability.
@@
where $\otimes$ denotes the outer product and $\mathbf{1}_{N}$ is a vector of ones. This operation effectively shifts the origin of the feature space to be class-specific, with the class token serving as the new zero point. The decoder then learns to reconstruct these recentered features rather than the original ones.
@@ 4.2 Comparison Results
(6) Few-Shot Multi-Class UAD. In resource-constrained scenarios where only limited normal samples are available, Dinomaly2 demonstrates exceptional few-shot learning capabilities. Using only 4 normal samples per class, our method achieves 98.1% I-AUROC on MVTec-AD and 96.7% on VisA, as shown in Table VII. These results even outperform the methods that were specially designed for few-shot anomaly detection, including IIPAD[74] (96.1%/88.3%) and WinCLIP[18] (94.0%/86.1%). Notably, Dinomaly2 surpasses SOTA full-shot MUAD models using only 8 samples per class.

### v2

@@ 1 Introduction
Early image UAD approaches follow a one-class-one-model paradigm[7, 8], where separate models are trained for each object category or scene (Fig. 1(a.1)). While this strategy has proven effective in controlled scenarios, it faces significant challenges in real-world applications. The storage overhead, computational complexity, and maintenance burden of managing hundreds of individual models become prohibitive in practical deployments. Recent advances have shifted toward multi-class UAD (MUAD, Fig. 1(a.2)), where a single unified model is designed to handle multiple object categories simultaneously[9, 10, 11]. Recent methods leverage a variety of techniques, including neighbor-masked attention [9], synthetic anomalies [10], vector quantization [12], diffusion model [13, 14], and state space model (Mamba) [15] to capture compact representations of normal patterns, which are critical for distinguishing anomalies. However, when intra-class normal patterns become excessively complex due to the presence of diverse categories, the underlying distribution becomes difficult to characterize, consequently degrading detection performance [9]. Yet despite this proliferation of complex designs, there is still a non-negligible performance gap between the state-of-the-art (SOTA) multi-class and single-class models (Fig. 1(d)), which restricts the practical adoption of unified approaches
@@
Moreover, we present simple strategies to extend vanilla 2D Dinomaly2 to achieve SOTA results on diverse scenarios including multi-view inspection, RGB+3D point cloud, RGB+Infrared (IR), and few-shot UAD (Fig. 1(b.2-4)). We further advance the field by introducing the inference-unified MUAD setting (Fig. 1(a.3)), where anomalies across mixed categories must be detected with a single threshold—a crucial step toward truly unified deployment. This setting exposes the limitations of existing methods that rely on category-specific calibration while demonstrating Dinomaly2’s robustness in this challenging setting.
@@ 2 Related Work
Unsupervised Anomaly Detection for computer vision has evolved through several paradigms, each addressing the fundamental challenge of identifying abnormal patterns without explicit anomaly supervision. Epistemic methods are based on the assumption that the networks respond differently during inference between seen input and unseen input [20]. Within this paradigm, pixel reconstruction [30] methods assume that the networks trained on normal images can reconstruct anomaly-free regions well, but poorly for anomalous regions. However, such models may also incorrectly reconstruct anomalous regions if these regions resemble normal regions[8]. Therefore, feature reconstruction (distillation) [8, 31, 32, 33] is proposed to construct features of pre-trained encoders instead of raw pixels. Recent advances [14, 34] have explored diffusion models for pixel and feature reconstruction. Synthetic anomaly methods generate pre-defined defects on normal images to imitate anomalies, converting UAD into supervised classification [35] or segmentation tasks [36, 37]. For example, CutPaste [35] simulates anomalous regions by randomly pasting cropped patches of normal images, while DRAEM [36] and DesTSeg [38] construct anomalous regions using Perlin noise as the mask and another image as the additive anomaly. SimpleNet [39] introduces anomaly by injecting Gaussian noise in the pre-trained feature space. These methods heavily rely on how well the synthetic anomalies match the real anomalies, which makes it hard to generalize to different datasets. Feature statistics (or memory-based) methods [40, 7, 41] memorize all normal features (or their modeled distribution) extracted by networks pre-trained on large-scale datasets and match them with test samples during inference, e.g., PatchCore[7]. Since these methods require memorizing, processing, and matching nearly all features from training samples, they are computationally expensive in both training and inference, especially when the training set is large.
@@ 3.4 Context-Aware Recentering
A fundamental challenge in multi-class UAD arises from the context-dependent nature of anomalies: identical visual features can be normal in one context but anomalous in another. For example, a vehicle is normal on a highway but anomalous on a pedestrian walkway, whereas a pedestrian is expected on a sidewalk but becomes anomalous on a highway (Fig. 4(a)). This contextual ambiguity becomes particularly problematic in unified multi-class models, where the decoder may learn to reconstruct both vehicles and pedestrians in any context (Fig. 4(b)), undermining the anomaly detection capability.
@@
where $\otimes$ denotes the outer product and $\mathbf{1}_{N}$ is a vector of ones. This operation effectively shifts the origin of the feature space to be class-specific, with the class token serving as the new zero point (Fig. 4(c)). The decoder then learns to reconstruct these recentered features rather than the original ones.
@@ 4.2 Comparison Results
(6) Few-Shot Multi-Class UAD. In resource-constrained scenarios where only limited normal samples are available, Dinomaly2 demonstrates exceptional few-shot learning capabilities. Using only 4 normal samples per class, our method achieves 98.1% I-AUROC on MVTec-AD and 96.7% on VisA, as shown in Table VII. These results even outperform the method that was specially designed for multi-class few-shot anomaly detection (IIPAD[73]: 96.1%, 88.3%). Moreover, our approach seamlessly integrates few-shot UAD with RGB-3D data. On the MVTec3D dataset, our method achieves an I-AUROC of 90.3% using only 4 samples per class, substantially outperforming the current SOTA[74] (74.0%)